can you help me with this equation?
one number is 15 more than another. The sum of twice the larger and 3 times the smaller is 182. Find the numbers.
Thank you

Answer by checkley77(12844)

y=x+15
2y+3x=182
2(x+15)+3x=182
2x+30+3x=182
5x=182-30
5x=152
x=152/5
x=30.4 ans. for the smaller number.
y=30.4+15
y=45.4 ans. for the larger number.
Proof:
2*45.4+3*30.4=182
90.8+91.2=182
182=182